EVA Report – January 2nd

EVA Report:

SOL: 01

Date: 02/01/2017

Person filling out the report: Pierrick Loyers

Crew members involved in the EVA: Ilaria Cinelli, Troy Cole, Patrick Gray, Gwendal Hénaff, Pierrick Loyers

EVA leader: Pierrick Loyers

Begin: 10.15 am

End: 12 pm

Type of EVA: Walking

Purpose: Preliminary tests of the scientific equipment

Coordinates: 38.245N 110.485W

Summary: Patrick, Gwendal and me have been the first of our crew to walk on Mars. Our objective was to transport outside the Ground Penetrating Radar which is pretty heavy. After waiting three minutes for depressurisation, we faced some difficulties to open the exit door. Several tries later, we finally manage to open the door and went outside where environmental conditions were pretty harsh due to strong wind. While Ilaria and Troy moved within the depressurisation sas to join us on the Martian soil, we turn on the GPR. Patrick was holding the 3D cartography system which well communicated with the Hab, transferring data through the radiofrequencies communication system. Anushree, who was in the Hab, kept an eye on the PC which was in communication with the cartography system outside, insuring that data were well received.

As soon as Ilaria and Troy finished their depressurisation time, we went all together to check diesel, propane and gasoline levels. Charging levels of Phobos and Diemos have been checked has well. When we moved to check water level, a bad surprise was waiting for us. Due to the strong wind, the protective cover on the water tank flew away and we caught it just at the right time. It took us approximately 20 minutes to put the protective cover back on the water tank and then we refilled the water tank.

After that, Ilaria, Gwendal, Troy and Patrick moved next to the ATVs in order to refill it on gasoline while I was preparing tests of the scientific equipment. Refilling ATVs has been difficult for the four crew members due to the wind.

Difficulties have been encountered with the GPR because due to my helmet, I was not able to visualize properly the screen and the measurements which has been an issue for the whole EVA. That was exactly the purpose of this EVA: to evaluate what could be the differences between wearing space suit or not while using the GPR. Despite of it, I have been helped by Gwendal and Patrick to drag the GPR all around the Hab. We have probe the soil next to the Green Hab and the solar panels because we suspected underground electrical pipes that could be detected thanks to the GPR. These pipes could also be used to calibrate the GPR.

While moving the radar around the Hab, the HUD that was carried by Gwendal ran out of battery and the 3D cartography system ran out of battery five minutes later as well. This is probably because the temperature was very cold this morning that batteries didn’t last a very long time. This is a problem in regards to the 3D cartography because data have to be collected on the device before it’s out of battery.

In total, 6 measurements have been made with the GPR and saved but the no data have been collected with the 3D cartography system. Despite of it, the long-range radiofrequency communication system worked well, transmitting data to the Hab. The HUD worked well as well while it has battery. The scientific equipment has been tested and so objectives of the EVA have been fully completed. Some negative points have been highlighted which was exactly the purpose of this EVA. These points will be improved for the incoming EVAs.